# On harmonic oscillator hazard functions

### v1

@@ 1 Introduction
The hazard function plays a key role in the analysis of survival data (Rinne, 2014). Given the intuitive interpretation of the hazard function as the instantaneous failure rate at time $t>0$, this function serves as basis for defining many survival regression models Rubio et al. (2019). Estimation of the hazard function using both parametric and non-parametric methods has received considerable attention. We refer the reader to Christen and Rubio (2024) for a comprehensive overview of these methods, which include the use of parametric distributions, splines, and Bayesian nonparametric methods. In the parametric setting, it is desirable to define models capable of capturing the basic shapes of the hazard function: increasing, decreasing, unimodal (up-then-down), and bathtub (down-then-up). To this end, several generalizations of the Weibull distribution have been proposed, including the generalized gamma distribution, the exponentiated Weibull distribution, and the power generalized Weibull distribution, among others. Similarly, general methods have been proposed to add parameters to a baseline distribution or baseline hazard function. See Marshall and Olkin (1997) and Anaya-Izquierdo et al. (2021) for an overview of these methods. More recently, Christen and Rubio (2024) proposed parametrically modeling the hazard function using a system of first-order autonomous ordinary differential equations (ODEs) with positive solutions. This approach offers a general methodology for generating distributions with interpretable parameters and allows for adding flexibility to the resulting solutions by including hidden states.
@@ 2 Modelling the hazard function with a harmonic oscillator
We propose modelling the hazard function, $h(t)$, through the linear second order ODE:
@@
Equation (1) models the acceleration of the hazard function and also has a geometric interpretation, describing the curvature of the hazard function, both at any time $t>0$. More specifically, this equation can be seen as a shifted version of the damped harmonic oscillator (Georgi, 1993; Strogatz, 2018). The damped harmonic oscillator is a classical model in Physics which is used to describe the motion of a mass attached to a spring when damping (friction) is present (Georgi, 1993; Strogatz, 2018). The damped harmonic oscillator is generally understood as a system that degrades or stabilizes over time due to the combined effects of the restoring force and damping. Thus, the damped harmonic oscillator (1) represents the evolution of the instantaneous failure rate, depicting a system that evolves over time to reach a degraded or stable state. Since the solutions to the damped harmonic oscillator equation can take negative values, and the equilibrium point is exactly 00, we consider the shifted version (1) of this model to allow for a positive equilibrium point $h_{b}>0$. The parameters of this ODE are easily interpreted. The natural frequency, $w_{0}>0$, controls the frequency of oscillation, while the damping coefficient, $\eta>0$, represents the dissipative forces acting on the system. The shift parameter, $h_{b}>0$, represents the stability state or equilibrium level of the solution. In the above parametrization the value of the damping coefficient $\eta$ determines the system’s behavior and, in our case, the shape of the hazard function. The three regimens of the system are the under-damped case ($10\leq\eta<$ 1), the over-damped case ($\eta>1$), and the critically damped case ($\eta=1$). We present a more detailed analysis of these cases in the next Sections.
@@ 2.4 Parametrization
Overall, we adopt the parametrization (1), with parameters $\eta$, $w_{0}$, $h_{b}$, $h_{0}$, and $r_{0}$. Nonetheless, other parametrizations might provide further insights into the model. Looking at the ODE formulation in 1, $w_{0}$ is called the “natural frequency” and its units are the inverse of time units. The parameter $\eta$ has no units, and is referred to as the “damping ratio”. Then, $w_{1}=w_{0}\sqrt{|\eta^{2}-1|}$ is the angular frequency in the under-damped case, and has the same units as $w_{0}$. This means that if $t\in[0,t_{M}]$, $t_{M}>0$, $h$ will oscillate $\frac{w_{1}t_{M}}{2\pi}$ times in that time window.
@@ 3 Shapes and tail-weight characterization
As previously discussed, one of the desirable properties of parametric survival models is their ability to capture the basic shapes. The following result presents a characterization of the hazard shapes of (1), in terms of the parameter values, which include the basic shapes as well as oscillatory cases. We focus on the under-damped and over-damped cases, which are the main cases of interest for modelling, as discussed in the previous Section.
@@
See Figure 1 for examples of the hazard shapes obtained as a solution to (1).
@@
Regarding the above result, an important difference of model (1) compared to the exponential distribution (which has a constant hazard function) or other parametric sub-exponential distributions (Vershynin, 2018), is that the proposed model (1) allows for capturing a variety of shapes of the hazard function. This sub-exponential tail-weight characterization also allows for deriving other properties of the distribution induced by (1), such as properties of the moments and moment generating function, as detailed in Chapter 2 of Vershynin (2018), and thus omitted here.
@@ 4 Inference
Since these functions are available in analytic form, the likelihood function can be implemented in any numeric programming language. This also implies that the maximum likelihood estimates can be calculated using general-purpose optimization methods.
@@ 6 Discussion
[…]

### v3

@@ 1 Introduction
The hazard function plays a key role in the analysis of survival data (Rinne, 2014). For a positive random variable $T$, with probability density function $f(t)$ and cumulative distribution $F(t)$, the hazard function is defined as:
$h(t)=\lim_{\Delta t\downarrow 0}\dfrac{{\mathbb{P}}\left(t\leq T<t+\Delta t\mid T\geq t\right)}{\Delta t}=\dfrac{f(t)}{1-F(t)}.$ (1)
Given the intuitive interpretation of the hazard function as the instantaneous failure rate at time $t>0$, this function serves as basis for defining many survival regression models Rubio et al. (2019). Estimation of the hazard function using both parametric and non-parametric methods has received considerable attention. We refer the reader to Christen and Rubio (2024) for a comprehensive overview of these methods, which include the use of parametric distributions, splines, and Bayesian nonparametric methods. In the parametric setting, it is desirable to define models capable of capturing the basic shapes of the hazard function: increasing, decreasing, unimodal (up-then-down), and bathtub (down-then-up). To this end, several generalizations of the Weibull distribution have been proposed, including the generalized gamma distribution, the exponentiated Weibull distribution, and the power generalized Weibull distribution, among others. Similarly, general methods have been proposed to add parameters to a baseline distribution or baseline hazard function. See Marshall and Olkin (1997) and Anaya-Izquierdo et al. (2021) for an overview of these methods. More recently, Christen and Rubio (2024) proposed parametrically modeling the hazard function using a system of first-order autonomous ordinary differential equations (ODEs) with positive solutions. This approach offers a general methodology for generating distributions with interpretable parameters and allows for adding flexibility to the resulting solutions by including hidden states.
@@ 2 Modelling the hazard function with a harmonic oscillator
[…]
@@
In any case, once $h$ is postulated as a hazard function, this fixes the distribution of the survival times, i.e. the probability model involved, and should always be interpreted as in (1). Equation (2) models the acceleration of the hazard function and also has a geometric interpretation, describing the curvature of the hazard function, both at any time $t>0$. More specifically, this equation can be seen as a shifted version of the damped harmonic oscillator (Georgi, 1993; Strogatz, 2018). The damped harmonic oscillator is a classical model in Physics which is used to describe the motion of a mass attached to a spring when damping (friction) is present (Georgi, 1993; Strogatz, 2018). The damped harmonic oscillator is generally understood as a system that degrades or stabilizes over time due to the combined effects of the restoring force and damping. Thus, the damped harmonic oscillator (2) represents the evolution of the instantaneous failure rate, depicting a system that evolves over time to reach a degraded or stable state. Since the solutions to the damped harmonic oscillator equation can take negative values, and the equilibrium point is exactly 00, we consider the shifted version (2) of this model to allow for a positive equilibrium point $h_{b}>0$. The parameters of this ODE are easily interpreted. The natural frequency, $w_{0}>0$, controls the frequency of oscillation, while the damping coefficient, $\eta>0$, represents the dissipative forces acting on the system. The shift parameter, $h_{b}>0$, represents the stability state or equilibrium level of the solution. In the above parametrization the value of the damping coefficient $\eta$ determines the system’s behavior and, in our case, the shape of the hazard function. The three regimens of the system are the under-damped case ($10\leq\eta<$ 1), the over-damped case ($\eta>1$), and the critically damped case ($\eta=1$). We present a more detailed analysis of these cases in the next Sections.
@@ 2.4 Parametrization
Overall, we adopt the parametrization (2), with parameters $\eta$, $w_{0}$, $h_{b}$, $h_{0}$, and $r_{0}$. Nonetheless, other parametrizations might provide further insights into the model. Looking at the ODE formulation in 2, $w_{0}$ is called the “natural frequency” and its units are the inverse of time units. The parameter $\eta$ has no units, and is referred to as the “damping ratio”. Then, $w_{1}=w_{0}\sqrt{|\eta^{2}-1|}$ is the angular frequency in the under-damped case, and has the same units as $w_{0}$. This means that if $t\in[0,t_{M}]$, $t_{M}>0$, $h$ will oscillate $\frac{w_{1}t_{M}}{2\pi}$ times in that time window.
@@ 3 Shapes and tail-weight characterization
As previously discussed, one of the desirable properties of parametric survival models is their ability to capture the basic shapes. The following result presents a characterization of the hazard shapes of (2), in terms of the parameter values, which include the basic shapes as well as oscillatory cases. We focus on the under-damped and over-damped cases, which are the main cases of interest for modelling, as discussed in the previous Section. See Figure 1 for examples of the hazard shapes obtained as a solution to (2).
@@
Although some parametric distributions, such as the power generalized Weibull or generalized gamma distributions, can capture the first four shapes, they do not account for oscillatory behaviors. Another important distinction is that the hazard function of the damped harmonic oscillator is always upper-bounded, with its amplitude decaying over time. Consequently, $h(t)$ approaches the stable state $h_{b}$ as $t\to\infty$. We acknowledge that these are implicit assumptions of the damped harmonic oscillator model, which may not be suitable for all types of data sets. However, we argue that all parametric models have implicit, and often overlooked, limitations. For instance, the Weibull distribution implicitly requires that when the hazard function is increasing, it must satisfy $h(0)=0$ and remain unbounded, which implies lighter tails than those of the exponential distribution (Christen and Rubio, 2024). Conversely, decreasing Weibull hazard functions must satisfy $h(0)=\infty$. The proposed damped harmonic oscillator model allows for a better understanding of the hazard shapes and properties thanks to the interpretability of the parameters in the ODE (2).
@@
The above result indicates that the distribution associated with the harmonic oscillator hazard function is a sub-exponential distribution (Vershynin, 2018). An important difference of model (2) compared to the exponential distribution (which has a constant hazard function) or other parametric sub-exponential distributions (Vershynin, 2018), is that the proposed model (2) allows for capturing a variety of shapes of the hazard function. This sub-exponential tail-weight characterization also allows for deriving other properties of the distribution induced by (2), such as properties of the moments and moment generating function, as detailed in Chapter 2 of Vershynin (2018), and thus omitted here.
@@ 4 Inference
Since these functions are available in analytic form, the likelihood function can be implemented in any numeric programming language. This also implies that the maximum likelihood estimates can be calculated using general-purpose optimization methods, although this may require comparing different methods an initial points since, in general, the concavity of the above likelihood (or log-likelihood) function is not guaranteed.
@@ 6 Discussion
The focus of this paper has been on survival analysis without covariates. A potential extension involves incorporating covariates into the damped harmonic oscillator hazard function proposed here. This can be achieved by using this model as the baseline hazard in any hazard-based regression model (Rubio et al., 2019), or by introducing covariates through linear predictors on the parameters (similar to distributional regression). These directions will be explored in future work.
@@
  • (i)
    In the under-damped case ($\eta<1$), we have that
    ${\mathbb{P}}(T>t\mid\bm{\theta})$ $=$ $\exp\left\{-H(t\mid\bm{\theta})\right\}$
    $=$ $\exp\left\{-h_{b}t\right\}\exp\left\{-\frac{A}{w_{0}\eta}\frac{\sin(\phi)+\mu\cos(\phi)}{\mu^{2}+1}\right\}$
    $\times$ $\exp\left\{\frac{A}{w_{0}\eta}\frac{e^{-w_{0}\eta t}(\sin(w_{1}t+\phi)+\mu\cos(w_{1}t+\phi))}{\mu^{2}+1}\right\}.$
    Now, using that $\frac{\sin(w_{1}t+\phi)+\mu\cos(w_{1}t+\phi)}{\mu^{2}+1}\leq\frac{1+\mu}{\mu^{2}+1}$, and denoting $K=\exp\left\{-\frac{A}{w_{0}\eta}\frac{\sin(\phi)+\mu\cos(\phi)}{\mu^{2}+1}\right\}$, $K_{1}=\dfrac{A}{w_{0}\eta}\dfrac{1+\mu}{1+\mu^{2}}$, and $K_{2}=w_{0}\eta$, it follows that
    ${\mathbb{P}}(T>t\mid\bm{\theta})\leq Ke^{-h_{b}t+K_{1}e^{-K_{2}t}}.$
    Therefore, we have shown that ${\mathbb{P}}(T>t\mid\bm{\theta})={\mathcal{O}}\left(e^{-h_{b}t+K_{1}e^{-K_{2}t}}\right)$.
  • (ii)
    In the over-damped case ($\eta>1$), we have that
    ${\mathbb{P}}(T>t\mid\bm{\theta})$ $=$ $\exp\left\{-H(t\mid\bm{\theta})\right\}$
    $=$ $\exp\left\{-h_{b}t\right\}\exp\left\{-\frac{h_{0}-h_{b}+a}{2(-w_{0}\eta+w_{1})}(e^{(-w_{0}\eta+w_{1})t}-1)\right\}$
    $\times$ $\exp\left\{-\frac{h_{0}-h_{b}-a}{2(-w_{0}\eta-w_{1})}(e^{(-w_{0}\eta-w_{1})t}-1)\right\}.$
    Now, denoting $K=\exp\left\{\frac{h_{0}-h_{b}+a}{2(-w_{0}\eta+w_{1})}\right\}\exp\left\{\frac{h_{0}-h_{b}-a}{2(-w_{0}\eta-w_{1})}\right\}$, $K_{1}=\max\left\{\dfrac{h_{b}-h_{0}-a}{w_{1}-w_{0}\eta},\dfrac{h_{0}-h_{b}-a}{w_{1}+w_{0}\eta}\right\}$, and $K_{2}=w_{0}\eta-w_{1}$, it follows that
    ${\mathbb{P}}(T>t\mid\bm{\theta})\leq Ke^{-h_{b}t+K_{1}e^{-K_{2}t}}.$
    Therefore, we have shown that ${\mathbb{P}}(T>t\mid\bm{\theta})={\mathcal{O}}\left(e^{-h_{b}t+K_{1}e^{-K_{2}t}}\right)$.
  • (iii)
    In the critically-damped case ($\eta=1$), we have that
    ${\mathbb{P}}(T>t\mid\bm{\theta})$ $=$ $\exp\left\{-H(t\mid\bm{\theta})\right\}$
    $=$ $\exp\left\{-h_{b}t\right\}\exp\left\{-\dfrac{(h_{0}-h_{b})\left(1-e^{-w_{0}t}\right)}{w_{0}}\right\}$
    $\times$ $\exp\left\{-\dfrac{\left(r_{0}+w_{0}(h_{0}-h_{b})\right)}{w_{0}^{2}}\left(e^{w_{0}t}-w_{0}t-1\right)e^{-w_{0}t}\right\}$
    Now, denoting $K=\exp\left\{-\dfrac{(h_{0}-h_{b})}{w_{0}}\right\}\exp\left\{-\dfrac{\left(r_{0}+w_{0}(h_{0}-h_{b})\right)}{w_{0}^{2}}\right\}$, $K_{1}=\max\left\{\dfrac{h_{0}-h_{b}}{w_{0}},\dfrac{r_{0}+w_{0}(h_{0}-h_{b})}{w_{0}^{2}}\right\}$, $K_{2}=w_{0}$, and $K_{3}=\dfrac{r_{0}+w_{0}(h_{0}-h_{b})}{w_{0}}$, it follows that
    ${\mathbb{P}}(T>t\mid\bm{\theta})\leq K\exp\left\{-h_{b}t\right\}e^{-h_{b}t+K_{1}e^{-K_{2}t}+K_{3}te^{-K_{2}t}}.$
    Therefore, we have shown that ${\mathbb{P}}(T>t\mid\bm{\theta})={\mathcal{O}}\left(e^{-h_{b}t+K_{1}e^{-K_{2}t}+K_{3}te^{-K_{2}t}}\right)$.
The result follows by noting that $K_{2}>0$ in all cases, which implies that the argument of the exponential can be upper-bounded by $-h_{b}t+C$, for some constant $C$, and $t$ sufficiently large. Consequently, the survival function has exponential tails ${\mathcal{O}}\left(e^{-h_{b}t}\right)$.